# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

### v2

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+23% on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@ 1 Introduction
Figure 1: Performance of Gemini Pro (v1.0) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023). The X-axis is the position of the relevant passage in the context. The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input. The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the main bottleneck is the identification of supporting evidence(s). DoubleDipper (our method, orange line) flattens this U-shaped trend.
@@
In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer. It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024). Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a primary performance bottleneck of current models in long contexts.
@@
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 23% on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
@@ 2 Background
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023). However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024). In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia. We describe the most relevant evaluation benchmarks used in this work in Section 4.
@@
While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context. The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024). In this work, we tackle these challenges and present a novel ICL method for long contexts.
@@ 3 DoubleDipper
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. By randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
@@ 4 Experiments
We apply our method to various datasets, each presenting its own domain-specific challenges. We selected these datasets because the supporting paragraphs are also annotated. Overall our evaluation set include 5.5K instances, with statistics of each dataset given in Table 1.
@@
We apply our method to a variety of models, both commercial and open-source. The commercial models include Gemini-1.0-Ultra and Gemini-1.0-Pro (Team et al., 2023). The open-source models we tested are Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Gemma-2b-it (v1.1), Gemma-7b-it (v1.1)(Team et al., 2024) and Mistral-7b-instruct (v0.2)(Jiang et al., 2023). Details about models’ size and context window are shown in Table 2.
For few-shot generation, we conducted two distinct experiments in order to analyze the effect of different question generation models. In the first experiment, dubbed DoubleDipper (Self), we employed the same LLM for both generating the demonstrations and answering to the query. In the second experiment, named DoubleDipper (PaLM 2), we generated a single set of few-shot examples using PaLM 2 (Anil et al., 2023) and then supplied these examples to various LLMs. This approach allowed us to assess the performance consistency across different models when provided with the same set of demonstrations.
@@
We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generate only the answer $a$.
Size Context Window
Gemini v1.0 Pro 32k
Ultra 32k
Llama-2 7B 4k
13B 4k
70B 4k
Mistral 7B 8k
Gemma 2B 8k
7B 8k
Table 2: Models used with DoubleDipper.
@@
We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate five QA pairs per passage (see Appendix A for the exact prompt). We then randomly select a single QA pair for each passage to form the few shot demonstrations. For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro 39.4 49.0 73.0 30.8 54.4 28.1 18.6 22.1
   + DoubleDipper (Self) 50.0 58.4 85.6 28.6 52.2 56.6 36.5 32.0
   + DoubleDipper (PaLM 2) 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra 53.9 53.0 81.2 30.8 66.0 61.6 42.6 42.3
   + DoubleDipper (Self) 53.1 57.2 73.4 20.6 54.6 68.9 51.5 45.6
   + DoubleDipper (PaLM 2) 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) 35.9 24.0 53.4 49.8 49.6 34.6 28.1 11.6
   + DoubleDipper (Self) 35.5 25.5 66.0 54.2 49.2 24.7 18.6 10.5
   + DoubleDipper (PaLM 2) 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) 27.2 7.1 52.2 49.8 48.6 12.3 16.7 3.4
   + DoubleDipper (Self) 47.6 45.8 92.6 74.2 45.4 40.4 22.4 12.4
   + DoubleDipper (PaLM 2) 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) 45.8 61.3 70.6 68.0 49.0 36.3 21.7 13.9
   + DoubleDipper (Self) 48.9 60.4 96.2 84.6 42.6 28.3 17.8 12.6
   + DoubleDipper (PaLM 2) 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat 31.6 42.0 46.8 39.0 37.4 32.7 15.1 8.2
   + DoubleDipper (Self) 33.3 45.6 64.2 37.2 35.4 23.3 17.8 9.2
   + DoubleDipper (PaLM 2) 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat 32.4 51.2 54.2 40.6 43.8 18.2 10.8 7.9
   + DoubleDipper (Self) 34.9 46.9 66.0 44.6 38.4 22.7 19.3 8.4
   + DoubleDipper (PaLM 2) 38.3 47.6 71.0 37.8 36.4 32.8 32.2 10.4
Llama-2-70b-chat 45.3 61.1 66.6 76.4 50.2 33.8 17.7 11.0
   + DoubleDipper (Self) 46.1 57.2 85.6 76.6 44.8 27.8 20.3 10.7
   + DoubleDipper (PaLM 2) 51.6 54.2 88.2 78.0 47.0 44.5 34.2 15.0
Table 3: Results of the baseline, DoubleDipper (Self) and DoubleDipper (PaLM 2) on various QA datasets.
@@ 5 Results
Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set. We report the performance of the supporting paragraphs prediction in Appendix B. On average across models and datasets, our straightforward method, which automatically generates few-shot examples from the provided input context, yields an improvement of 12% for DoubleDipper (Self) and 23% for DoubleDipper (PaLM 2), often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B). Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).
Furthermore, following (Liu et al., 2023), Figure 4 shows the performance of the tested LLMs for both the baseline and DoubleDipper (Self) on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. For Gemini Pro, Gemini Ultra, Gemma 2B, Gemma 7B, Llama 7B, the performance curve for DoubleDipper consistently surpasses the baseline across nearly all document positions. This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document. For Mistral, Llama 2 13B and Llama 2 70B, Figure 4 shows that DoubleDipper performs on par or better than the baseline when the relevant document is in the middle of the input, but still underperforms the baseline at the beginning and the end. This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias.
@@
Figure 4: Performance (accuracy) of the models on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
6 Analysis
Few-shot vs. Zero-shot
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Zero-shot 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra Zero-shot 56.5 56.2 76.8 34.8 62.2 68.3 52.4 44.9
DoubleDipper 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) Zero-shot 3.6 6.4 4.2 5.4 0.0 3.3 4.9 0.7
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) Zero-shot 7.5 2.6 36.0 5.4 0.0 3.9 3.3 1.0
DoubleDipper 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) Zero-shot 41.9 59.0 77.4 67.4 51.8 13.7 12.3 11.8
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat Zero-shot 22.4 29.2 45.0 29.2 33.4 9.8 7.9 2.4
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat Zero-shot 14.0 8.0 43.0 3.0 37.0 5.2 0.3 1.2
DoubleDipper 38.0 36.7 71.4 36.6 50.0 27.4 27.4 16.8
Table 4: Comparison of DoubleDipper to prompting models in a zero-shot setting to provide explicitly the supporting paragraphs, then generate the answer.
In order to assess the effectiveness of the few shots in DoubleDipper, we prompt the model in a zero-shot setting to explicitly specify the ID of the relevant passage(s) before generating the answer. The results of the QA task are detailed in Table 4 and the performance in identifying supporting paragraphs is reported in Appendix B.
For all open-source models, this zero-shot prompting strategy leads to a significant decline in QA performance compared to the baseline (see Table 3), which merely prompts the model for the answer (e.g., -32.3 points for Gemma 2B, -9.2 points for Llama 7B, -4.2 points for Mistral). This drop is likely because retrieving relevant passages makes the task more complex compared to standard QA, as models are not typically trained to perform such “retrieval” tasks. This is further supported by the poor performance in supporting passages identification shown in Appendix B, where F1 scores are often close to zero.
Despite these challenges, DoubleDipper’s self-generated demonstrations systematically improve performance for all open-source models on both QA and supporting passages identification across all datasets. On average, Mistral’s QA performance improves by 10.9 points, and its identification of relevant paragraphs increased by 30.5 F1 score. This aligns with previous research suggesting that models can learn new tasks through in-context learning (ICL).
Comparison to traditional ICL
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro ICL 48.6 59.1 69.8 30.0 52.8 55.6 41.0 32.0
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemma-2b-it (v1.1) ICL 21.6 0.2 28.2 26.4 13.0 39.7 39.1 4.3
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Llama-2-7b-chat ICL 16.7 0 13.2 13.2 14.2 37.0 38.2 0.7
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Mistral-7b-instruct ICL 52.1 68.5 79 67.8 39.2 49.4 36.7 24.1
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Table 5: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
To analyze the performance gain of DoubleDipper compared to the traditional In-Context-Learning, we prompt a representative model from each family (Gemini Pro, Gemma 2B, Llama 7B, and Mistral) with few-shot examples from the same dataset, each example including its own context and question-answer pair.
We report the results in Table 5. For all models, DoubleDipper achieves a higher performance than the traditional ICL. Interestingly, for Gemma 2B and Llama 7B, the traditional ICL not only underperforms DoubleDipper but it also significantly hurts performance compared to the baseline (e.g., -14.3 for Gemma 2B and -14.9 for Llama 7b). This performance drop might be because each demonstration in the traditional ICL setup comprises a few thousand tokens, which exacerbates the models’ existing challenges with processing long-range dependencies.

### v3

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@ 1 Introduction
Figure 1: Performance of Gemini Flash (v1.5) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023). The X-axis is the position of the relevant passage in the context. The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input. The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the identification of supporting evidence(s) is a major challenge. DoubleDipper (our method, orange line) flattens this U-shaped trend.
@@
In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer. It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024). Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a major performance bottleneck of current models in long contexts.
@@
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro, Nano and Flash; Reid et al., 2024) and open-source (Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Gemma (Riviere et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 16 absolute points on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
@@ 2 Background
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Hsieh et al., 2024); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023). However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024). In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia. The problem of designing informative and reliable benchmarks in long-context is an an active, ever-changing area of research (Goldman et al., 2024; Yen et al., 2024). We describe the most relevant evaluation benchmarks used in this work in Section 1.
@@
While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context. The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024c).
More related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks (Kim et al., 2022; Yasunaga et al., 2024; Li et al., 2024a, b). However, applying those methods to long contexts is not straightforward because LLMs would need to generate long demonstrations.
In this work, we tackle these challenges and present a novel ICL method that recycles the given long context and further instructs models to identify the relevant information before generating the answer.
@@ 3 DoubleDipper
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. Also, by randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation compared to the traditional In-Context-Learning. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
@@ 4 Experiments
Size Context Window
Gemini v1.5 Nano 32k
Flash 32k
Pro 32k
Llama 3.1 7B 128k
Mistral v0.3 7B 32k
Mixtral v0.1 7x8B 32k
Gemma 2 2B 8k
9B 8k
Table 1: Models used with DoubleDipper.
@@
We apply our method to various datasets, each presenting its own domain-specific challenges. We selected these datasets because the supporting paragraphs are also annotated. Overall our evaluation set includes 5.5K instances, with statistics of each dataset given in Table 2.
@@
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemini Pro (vanilla) 60.5 24.9 95.0 97.6 64.4 46.7 71.6 23.1
   Zero-shot + Evidence Retrieval 62.3 32.5 94.6 95.8 62.4 49.6 74.8 26.5
   DoubleDipper 70.4 46.8 97.4 99.0 79.6 60.9 72.4 36.4
Gemini Flash (vanilla) 42.9 10.2 70.0 86.0 57.6 10.0 59.5 7.3
   Zero-shot + Evidence Retrieval 58.2 30.2 78.8 90.6 65.0 44.9 67.4 30.2
   DoubleDipper 66.1 48.0 85.8 95.0 68.6 60.6 65.0 39.7
Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
Gemma 2 9B (v2) (vanilla) 44.0 11.4 74.8 81.8 55.6 13.8 61.0 9.3
   Zero-shot + Evidence Retrieval 58.7 38.6 82.0 83.4 59.4 56.1 64.4 26.8
   DoubleDipper 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
Llama 3.1 8B (vanilla) 37.2 11.8 56.2 52.2 48.6 20.7 63.5 7.4
   Zero-shot + Evidence Retrieval 53.1 42.2 71.8 65.4 50.8 55.3 62.0 24.5
   DoubleDipper 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
Mistral 7B (v0.3) (vanilla) 37.4 14.1 59.2 57.4 50.2 15.8 60.8 4.6
   Zero-shot + Evidence Retrieval 44.0 23.8 66.2 62.4 49.6 34.1 58.6 13.6
   DoubleDipper 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
Table 3: Accuracy of the QA task for the vanilla baseline (prompting the model to only answer the question), Zero-shot + Evidence Retrieval (prompting the model to explicitly identify the relevant passage(s) before generating the answer) and DoubleDipper with 3 demonstrations generated by Gemma 2 2B.
@@
We apply DoubleDipper to a variety of models, both commercial and open-source. The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash and Gemini 1.0 Nano (Reid et al., 2024). The open-source models we tested are Llama 3.1 8B (Dubey et al., 2024), Gemma 2B (v2) and Gemma 9B (v2) (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) and Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024). Details about models’ size and context window are shown in Table 1.
Few-shot generation in DoubleDipper is an auxiliary task and should ideally run in an efficient time without requiring heavy resources. Therefore, in our main experiments, we employ Gemma 2B to generate the demonstrations at it is the smallest and most efficient model used in our experiments. See Section 6 for an ablation analysis of the effect of the chosen model for generating the demonstrations.
@@
We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generates only the answer $a$. We also compare to Zero-shot + Evidence Retrieval, which prompts the model in a zero-shot setting to identify the relevant passage(s) before generating the answer.
@@
We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate a single QA pair for each passage (see Appendix A for the exact prompt). For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
@@ 5 Results
Result 1: DoubleDipper offers a substantial performance boost.
Table 3 presents the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. The results first show that prompting models to explicitly identify the relevant paragraphs before generating the answer (Zero-shot + Evidence Retrieval) leads to a performance improvement of 9.7 points on average across models over the vanilla baseline. DoubleDipper offers an additional substantial boost of 6.3 points for all models on average, culminating in a overall improvement of 16 absolute points over the vanilla baseline. Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique). Likewise, DoubleDipper outperforms the baseline also on the FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemini Pro Zero-shot + Evidence Retrieval 83.7 96.7 97.7 97.5 62.8 92.1 63.6 75.3
DoubleDipper 86.3 94.4 99.8 97.1 80.9 90.0 66.4 75.4
Gemini Flash Zero-shot + Evidence Retrieval 75.7 82.3 90.5 72.6 70.4 80.9 67.3 65.9
DoubleDipper 79.1 83.7 98.3 80.2 71.2 84.6 66.1 69.5
Gemini Nano Zero-shot + Evidence Retrieval 68.1 76.3 77.6 71.6 66.0 74.1 55.8 55.2
DoubleDipper 69.9 76.6 86.6 74.5 59.8 76.0 59.2 56.7
Gemma 2B Zero-shot + Evidence Retrieval 39.1 57.6 49.4 44.3 18.9 53.1 14.0 36.4
DoubleDipper 45.6 56.5 60.7 57.5 13.1 58.1 33.1 39.9
Gemma 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
DoubleDipper 57.0 74.2 52.5 59.0 23.1 78.4 55.3 56.8
Llama 3.1 8B Zero-shot + Evidence Retrieval 61.7 53.2 86.5 66.9 67.8 63.2 41.1 53.5
DoubleDipper 61.4 68.9 71.4 54.9 52.8 73.7 53.0 54.8
Mistral 7B (v0.3) Zero-shot + Evidence Retrieval 46.6 62.4 49.8 46.2 17.5 64.0 43.4 43.0
DoubleDipper 43.8 63.4 33.8 42.5 4.2 66.7 55.6 40.2
Mixtral 7x8B v(0.1) Zero-shot + Evidence Retrieval 60.0 72.4 69.4 64.6 43.4 76.9 40.6 52.4
DoubleDipper 58.9 70.4 81.6 63.6 18.3 75.0 50.4 53.2
Table 4: Performance (F1) of supporting paragraph(s) prediction.
@@
Figure 4: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial models than in open source.
Table 4 presents the performance of the supporting paragraphs prediction for the Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. For all commercial models and Gemma 2B, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +1.8 F1 for Gemini Nano and +6.5 F1 for Gemma 2B). On the other hand, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral and -0.3 F1 for Llama 3.1). This difference is because DoubleDipper’s demonstrations are based on a single paragraph, where commercial models can generalize better and predict multiple evidences. Indeed, Gemini Pro predicts on average 2 evidences for the datasets 2Wiki, MuSique, PIR, MonoRel, SRT and HotPotQA whereas Gemma 9B predicts 1.2 evidences for each instance. In fact, the only dataset with a single evidence is “Lost” and almost all models highly benefit from DoubleDipper (e.g, +11.9 F1 for Llama 3.1, +12.2 F1 for Mistral 7B and +9.8 for Mixtral 7x8B).
Result 3: DoubleDipper makes models more robust to the position of relevant information.
Following Liu et al. (2023), Figure 4 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. See Appendix B for the performance curve of the other tested models, which show similar trends.
Overall, the performance curve for DoubleDipper consistently surpasses the baseline when the relevant information appears “in the middle” and sometimes also at the beginning and/or the end (e.g., Gemini Flash). This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias. This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document.
6 Ablation Studies
Gemma 2B Self Gemini Pro
Gemini Pro 70.4 71.6 71.6
Gemini Nano 62.1 61.7 62.8
Gemini Flash 66.1 67.5 68.1
Gemma 2B 49.5 49.5 51.0
Gemma 9B 61.2 62.0 63.5
Llama 3.1 59.9 60.1 61.5
Mistral v0.3 51.0 49.8 52.6
Mixtral 52.2 49.8 54.4
Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix C.2 for the results on each evaluation dataset.
How many examples are needed?
In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10. On average, a single demonstration already provides an improvement over the baseline. Adding 3 demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b). We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
$k=1$ $k=3$ $k=5$ $k=10$
Gemini Nano 60.0 62.1 62.2 62.3
Gemini Flash 65.3 66.1 65.9 66.1
Gemma 2B (v2) 47.0 49.5 49.6 49.9
Gemma 9B (v2) 58.7 61.2 61.4 61.3
Llama 3.1 57.7 59.9 60.6 61.4
Mistral 7B (v0.3) 48.9 51.0 51.1 51.4
Mixtral 7x8B (v0.1) 49.3 52.2 51.7 52.2
Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix C.1 for the results on each evaluation dataset.
Investigating the effect of the few-shot generator
To understand the impact of the default chosen model (Gemma 2 2B) for generating the demonstrations, we conducted two additional experiments. The first experiment is Self in which we use the same model for generating the demonstrations and for answering the original question. In the second experiment, we generate the demonstrations with the best LLM used in our experiments, namely Gemini Pro. The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix C.2. The results show that generating the demonstrations with Gemma 2 or Self achieves similar performance, while Gemini Pro leads to a consistent increase in performance across models, indicating that future better models can improve further the performance.